You have a stock 4.00 times 10^-4 solution of KMnO
Answer #1

1.2 0.8 0.4 0.2 0.0001 0.0002 0.0003 0.0004 0.0005 Concentration

The relevent equation is y = 2637.5x + 0.022 where y = abs and x = conc.

unknown concentration A is = (0.83-0.022)/2637.5 = 3.06 *10^-4 M

unknown concentration B is = (0.34-0.022)/2637.5 = 1.21 *10^-4 M

Volumn, 하 Ab8 stoek ute 40 20 1O comen 1o5 4X16 M 45 O 22 4 x 104 x 40 50 3.2 X10 M 24x 04x 10 50 Extinetion Coetticienmol J